OpenStudy (tokyoqu33n):

NEED HELP ASAP Which ordered pairs are on the line with equation y+3=2(x−2) Select each correct answer. (3.5, 0) (2, -3) (4, 1) (0, -2)

OpenStudy (tkhunny):

Please substitute the ordered pairs into the equation. Do they make sense?

OpenStudy (tokyoqu33n):

?

OpenStudy (tkhunny):

(1,2) MEANS x = 1 and y = 2

OpenStudy (tokyoqu33n):

i know that but it ask for more than one

OpenStudy (tkhunny):

Then what's stopping you? Try all four. See if they make sense in the equation.

OpenStudy (tokyoqu33n):

uuuuuuugh god nvm

OpenStudy (tkhunny):

Why don't you do it? y+3=2(x−2) (3.5,0) So, 0 + 3 = 2(3.5-2) Then 3 = 2(1.5) Then 3 = 3 Well, that makes sense, so (3.5,0) is good. You do the next one.

OpenStudy (tokyoqu33n):

i said nvm cause my brother is helping me now

OpenStudy (tokyoqu33n):

thx anyway .....ig

OpenStudy (tkhunny):

We like helpful brothers. Hang in there. You'll get it.

OpenStudy (tokyoqu33n):

ig so i want him to show me how to get the answer but hes telling me the answer's cause we have to go